### v3

@@ 4.1.2 Updates on forward modelling of slow rotators
The Ledoux rotational splitting in Eq. (14) is a good approximation for the p modes of low-mass dwarfs, for the p modes of subgiants and red giants, for the p and g modes of subdwarfs, and for the g modes of white dwarfs. For most of these pulsators, the oscillation periods range from several minutes to a few hours and are small fractions of the rotation periods ranging from about a day for white dwarfs (Hermes et al. 2017) to several months for red giants (Mosser et al. 2012; Gehan et al. 2018). Aside from a few exceptions, the spin parameters of the modes of all these types of pulsators in the co-rotating frame, , put them far into the super-inertial regime (completely to the right in Fig. 3). Almost all of these pulsators are hence slow rotators in our definition, with their forward modelling being done from their mode frequencies and the subsequent derivation of from Eq. (14) relying on the best forward seismic model. With the notable exception of the high-order p modes in the Sun-like star Bootis modelled with inclusion of the centrifugal deformation by Suárez et al. (2010), summaries of asteroseismic modelling results for the slow rotators can be found in García & Ballot (2019), Hekker & Christensen-Dalsgaard (2017), Lynas-Gray (2021), and Giammichele et al. (2022) for low-mass dwarfs, red giants, subdwarfs, and white dwarfs, respectively. Homogeneous analyses for the subgiant phase were somewhat lacking after the Kepler mission finished, but are now also well underway (Ong et al. 2021; Noll et al. 2021), in anticipation of TESS light curves for this evolutionary phase.
@@ 4.1.3 Detection of internal rotation and core magnetism in red giants
The CoRoT space mission revealed red giants to be non-radial pulsators (De Ridder et al. 2009). It was a lucky circumstance that the mission program could not avoid red giants in the observing fields dedicated to exoplanet research. The mode periods roughly cover from half an hour to half a day, while the rotation periods of red giants typically cover from ten to hundreds of days. As discussed in the previous section, this implies that rotation can be treated perturbatively following the Ledoux approximation and modelling can be done from the zonal modes. The CoRoT discovery of non-radial oscillations in red giants allowed the application of age-dating for various areas of the Milky Way (Hekker et al. 2009; Miglio et al. 2009) with major advances for galactic archaeology (e.g., Miglio et al. 2013; Chiappini et al. 2015; Montalbán et al. 2021) as one of the important spin-offs of asteroseismology as indicated in Fig. 1.
@@
Following the theoretical predictions from Dupret et al. (2009) triggered by CoRoT, dipole mixed modes were discovered in CoRoT data of a subgiant (Deheuvels et al. 2010) and in Kepler data of a red giant (Beck et al. 2011). Subsequently, such modes were found in a whole sample of red giants by Bedding et al. (2011). The latter breakthrough study led to the important capacity to discriminate between stars on the red giant branch (RGB) and in the first or secondary red clump, even though such stars share the same surface properties. This discriminating capacity was also nicely illustrated for red giants in open clusters (Stello et al. 2011) and relies on the fact that the mixed modes have g-mode character in the deep interior of the star, yet are of acoustic nature in the envelope. Looking at Fig. 3, one can understand the nature of these modes by realising that the p- and g-mode cavities delineated by the and symbols come closer to each other as a star evolves from being a dwarf to a red giant. As shown by Dupret et al. (2009), this narrows the zone between and in such a way that the waves can tunnel through both cavities without becoming fully evanescent in the transition zone, thus creating dipole mixed modes that probe the entire star.
@@
Core rotation frequencies were detected from the splitting of mixed dipole modes in a few Kepler red giants after two years of uninterrupted photometric monitoring (Beck et al. 2012). The positive effect of the longer duration of the Kepler light curves beyond two years implied major progress in the seismic precision (Hekker et al. 2012). It led to a revolution in the understanding of angular momentum transport from large samples of evolved stars initiated by Mosser et al. (2012), with refined analyses from rotation inversions for a few RGB pulsators by Deheuvels et al. (2012) and red clump giants by Deheuvels et al. (2015). Following the initial breakthroughs, the internal rotation of evolved stars became an industrialised observational science of major importance (e.g., Gehan et al. 2018).
@@
Space asteroseismology meanwhile delivered the internal rotation for several thousands of evolved low- or intermediate-mass stars, notably subgiants and their successors climbing up the red giant branch, as well as giants in the first and secondary red clumps, subdwarfs, and white dwarfs. A summary of the rotational properties of about 1200 slowly rotating field stars is available in the review paper by Aerts et al. (2019) and not repeated here. Fig. 4 in that paper, along with the updated figures with measured internal rotation rates until end 2019 in Aerts (2021), place the internal rotational properties of low- and intermediate-mass stars in a global evolutionary picture. This revealed that the angular momentum of the core of helium burning red giants is in agreement with the angular momentum of white dwarfs.
@@
The major new insight from asteroseismology regarding internal rotation across stellar evolution is that the level of non-rigidity in the internal rotation of single dwarfs of low and intermediate mass is small during their longest phases of evolution, when they fuse hydrogen into helium and helium into carbon/oxygen. As discussed in Aerts et al. (2019), who summarised all results obtained until mid-2018, any explanation of this asteroseismic picture of stellar evolution requires more efficient transport of angular momentum during these phases than anticipated prior to Kepler space asteroseismology. Said differently, much stronger coupling between the convective core and envelope occurs during the two central burning stages of intermediate-mass stars. New theories to address this too low angular momentum transport in models have meanwhile been developed. One way to increase the angular momentum transport while stars evolve is to include internal gravity waves in the models. The occurrence of such waves was already proposed as an efficient way to transport angular momentum in low-mass stars like the Sun by Charbonnel & Talon (2005); Rogers et al. (2013) and in high-mass stars by Rogers et al. (2013).
@@
Another way to achieve higher angular momentum transport in models is by including magnetic effects. The existence of internal magnetic fields in red giants was already inferred by Fuller et al. (2015) as an interpretation for depressed dipole mixed modes found in Kepler data for a small fraction of red giants. The accompanying study by Cantiello et al. (2016) showed that a so-called magnetic greenhouse effect may be operational in red giants, turning g modes trapped in the core into Alfvén waves. The mode damping requires internal magnetic field strengths above about G, as confirmed by the theoretical predictions in Loi & Papaloizou (2017). A test of this greenhouse scenario for earlier phases of stellar evolution was developed by Stello et al. (2016), namely about half of the F dwarfs should have a strong internal magnetic field during their main-sequence phase. So far, it has not been possible to test this hypothesis from a representative asteroseismic population of F dwarfs, but internal magnetic field strengths of the 37 best modelled Kepler Dor pulsators inferred by Aerts et al. (2021) are in agreement with the scenario by Fuller et al. (2015) keeping in mind the minimal field strengths deduced by Cantiello et al. (2016).
@@
Although the interpretation by Fuller et al. (2015) was contested by Mosser et al. (2017), the presence of internal magnetic fields in a fraction of red giants is no longer a hypothesis. Indeed, an observational breakthrough was achieved by Li et al. (2022a) and by Deheuvels et al. (2023), who found evidence for the presence of such fields inside the cores of three, respectively eleven red giants from 4-yr Kepler light curves. Li et al. (2022a) detected the core magnetism from the occurrence of asymmetrical rotational splittings in the triplets of mixed dipole modes. They interpreted this asymmetry as being due to the small effects induced by the Lorentz and Coriolis forces on the modes. In their approximation, the terms caused by and the magnetic force in are considered to be small compared to those due to the gas pressure gradient and gravity in Eq. (1), as confirmed by their numerical computations. Elegant analytical and numerical work to interpret the observational findings in the adopted approximation was offered by Li et al. (2022a, see the Supplementary Material) and by Mathis & Bugnet (2023) for various magnetic field topologies, following earlier theoretical studies on the effect of magnetism on modes of red giants by Loi & Papaloizou (2018); Gomes & Lopes (2020); Loi & Papaloizou (2020); Loi (2020b, a, 2021); Bugnet et al. (2021); Bugnet (2022). Deheuvels et al. (2023), on the other hand, inferred the presence of a core magnetic field in eleven RGB stars from the effect of the Lorentz force on the frequencies of zonal dipole modes. These eleven red giants do not reveal asymmetries in their dipole mode splittings. Rather, the Lorentz force affects the dipole zonal mode frequencies of consecutive radial order. The effect depends on the strength of the magnetic field and allows to deduce lower limits on the core field strengths. As such, Deheuvels et al. (2023) found fields of 40 kG to 610 kG for these eleven stars. One of them, KIC 6975038, has depressed dipole modes, lending further support to the original interpretation of a magnetic greenhouse effect by Fuller et al. (2015).
@@
Meanwhile Li et al. (2023a) detected a core magnetic field from asymmetrical splittings in thirteen of the red giants with detected dipole mixed mode triplets found in the sample of 8000 compiled by Yu et al. (2018) and Gehan et al. (2018). The way the authors selected these 1200 RGB stars disfavours such stars with depressed dipole modes, as well as stars in an advanced stage of evolution along the RGB. The average core magnetic field strengths of these thirteen stars range from 20 to 150 kG, representing a range from 5% to 30% of the so-called critical field strength above which magneto-gravity waves get dissipated in the core rather than being propagative (Fuller et al. 2015; Rui & Fuller 2023). The detected fields were found to have a variety of horizontal field geometries. The core magnetic fields detected in less than 2% of the Kepler red giants with dipole mixed modes by Li et al. (2023a) and Deheuvels et al. (2023) are important to improve stellar evolution models. The low fraction found in the large homogeneous database of the Kepler 4-yr light curves does not necessarily imply the absence of a field for the majority of stars as the fields may remain undetected for large parts of the evolution. Indeed, the field strengths were found to decrease as the cores of the stars shrink along their evolutionary path. The magnetic red giants have core rotation properties fully in line with those of the thousands of red giants without detected magnetic signatures. Both findings suggest that the magnetic fields do not cause much extra transport of angular momentum compared to the case of a non-detectable magnetic field. The sample of red giants with detected core fields is currently small to make strong general conclusions. The effects of the core fields on internal mixing, if any, remain to be studied given the recent character of these discoveries. This may help to assess if the magnetic field has any effect on element transport as the red giants evolve.
@@ 4.2 Asteroseismology of supernova progenitors: low-order p and g modes in Cep stars
Moving somewhat to the left on the frequency axis in Fig. 3, we find moderate rotators among the Cep stars, such as HD 192575 already discussed above. The Cep stars are dwarfs or (super)giants of spectral type O or early B with masses roughly in the range 8 to 25 M. They are hence precursors of supernovae. This class of pulsators has always played a special role in asteroseismology, and this is no different in this review. As early as 1902, Frost (1902) discovered radial-velocity variations for the prototype of the class, the star Cephei. At that time, the only explanation was that the star is a spectroscopic binary. Although the variability of tens of similar stars was misinterpreted in terms of orbital motion for about half a century, a breakthrough was achieved thanks to the three periodicities found in the radial-velocity time series for the class member CMa (Struve 1950; Van Hoof & Struve 1953). These properties in fact led Ledoux (1951) to write his breakthrough paper on non-radial oscillation modes in a rotating star, introducing the concept of rotationally split multiplets due to the joint effect of the Coriolis force and the projection in the line-of-sight for an observer, as discussed in Sect. 3.3. Even though we know today that Ledoux’s interpretation of the variability of CMa was wrong (e.g. Mazumdar et al. 2006, for a modern asteroseismic interpretation), it led him to derive a general description of the observed characteristics of non-radial oscillations in a rotating star, which is the one still in use today for slow rotators with modes in the super-inertial regime.
@@
Following Goupil et al. (2000), we assess the relative importance of 2nd-order effects due to the Coriolis force, which scales as , and the one due to the centrifugal force, which is proportional to . Thus, the relative role of the centrifugal versus Coriolis forces in any 2nd-order theory in behaves as
@@
This ratio shows that the Coriolis force dominates over the centrifugal force for high-order g modes in many of the brightest SPB pulsators, for which this ratio has been measured and is typically between 0.01 and 0.1 (see Fig. 17 in De Cat & Aerts 2002). One can therefore ignore the flattening caused by the centrifugal force for such moderately rotating pulsators, but one cannot ignore the Coriolis force in their asteroseismic modelling (Aerts et al. 2021). Ballot et al. (2010) unravelled the domains of validity of a perturbative approach for the rotation and concluded that the 2nd-order approach still gives satisfactory results for equatorial velocities up to 100 km s for Dor stars and up to 150 km s for SPB stars. A 3rd-order perturbative approach increases the domains of validity by a few tens of km s. However, we know since the space asteroseismology era that most of these pulsators have their modes in the sub-inertial regime (see Fig. 3 and Aerts et al. 2021) where perturbative approaches for the rotation are not valid. We treat such high-order g-mode pulsators by adopting the so-called Traditional Approximation of Rotation (TAR) in Sect. 4.4, where it will be explained that the TAR is a mathematically elegant approximation that ignores the horizontal component of the rotation vector.
@@
The ratios from Eq. (15) for typical p modes in Cep or Sct stars are roughly between 1 and 10 (Goupil et al. 2000). This explains why the low-order modes in slowly to moderately rotating Cep and Sct stars can still be described perturbatively by a 2nd-order perturbative theory while ignoring the flattening due to the centrifugal force in the equilibrium models. Ballot et al. (2010) evaluated the validity domain for these p-mode pulsators and concluded that this approximation is fine for equatorial rotation velocities up to about 70 km s. They stressed that adding 3rd-order terms hardly improves this because p modes are only weakly sensitive to the Coriolis force. For the cases for which low-order linear oscillation modes in moderately rotating Cep stars can still be described perturbatively, we refer to Soufi et al. (1998) and Karami (2008) for appropriate mode expressions.
@@
Rotationally split p and g modes should in principle be easy to recognise for the slower rotators in the class and help identify the modes. Yet, despite the hundreds of OB pulsators found in space photometry (Aerts et al. 2006a; Balona et al. 2011; Balona 2016; Balona et al. 2019; Pedersen et al. 2019; Burssens et al. 2019, 2020; Balona 2022), the frequency spectra turn out to contain so many frequencies that the modes are hard to identify from the single-passband photometric variations. In particular, zonal modes are hard to find. As a result, forward modelling in the space era has been done for only few of these supernova progenitors (Aerts et al. 2006b; Handler et al. 2009; Briquet et al. 2011; Aerts et al. 2011; Daszyńska-Daszkiewicz et al. 2017; Handler et al. 2019). Given their mass range and extensive convective cores, the forward modelling problem to fit the zonal modes is at minimum 5-dimensional, because convective core overshooting cannot be ignored and is an uncalibrated mixing phenomenon requiring at least one fit parameter.
@@
Some of the most detailed forward models for this class result from earlier months-long ground-based multi-site campaigns for several bright class members (Pamyatnykh et al. 2004; Briquet et al. 2007; Dziembowski & Pamyatnykh 2008; Briquet et al. 2012). A detailed comparison among the forward models resulting for the three pulsators HD 129929, Oph and V2052 Oph, which have very similar dominant modes including an quintuplet, a radial mode, and an triplet, revealed that the magnetic star V2052 Oph has lower core overshooting than the two other non-magnetic pulsators. This hinted at inhibition of mixing by a magnetic field.
@@
So far, internal rotation profiles assuming shellular rotation have been determined for five Cep stars. This had led to ratios of the near-core and envelope rotation rates between 1 and 5 (cf. Table 1 and Fig. 5 in Burssens et al. 2023, for a summary), where the rigid rotation case occurs in the only close binary pulsator among this small sample. However, the methods adopted to derive the internal rotation are diverse. For the extensively studied star Eri, Dziembowski & Pamyatnykh (2008) adopted the 1st-order Ledoux approximation according to Eq. (14) and used the fundamental radial mode and the two triplets and to obtain a core-to-envelop rotation ratio of about five. Suárez et al. (2009), however, performed a more detailed analysis based on a 2nd-order perturbative approach, including as well the triplet. They found a factor lower envelope than near-core rotation from the three observed triplets, while also being able to explain most (but not all) of the measured triplet asymmetries. The approach by Suárez et al. (2009) is a promising route to understand angular momentum redistribution and chemical mixing due to differential rotation and its induced turbulence in moderately rotating Cep stars.

### v4

@@ 4.1.2 Updates on forward modelling of slow rotators
The Ledoux rotational splitting in Eq. (14) is a good approximation for the p modes of low-mass dwarfs, for the p modes of sub-giants and red giants, for the p and g modes of sub-dwarfs, and for the g modes of white dwarfs. For most of these pulsators, the oscillation periods range from several minutes to a few hours and are small fractions of the rotation periods ranging from about a day for white dwarfs (Hermes et al. 2017) to several months for red giants (Mosser et al. 2012; Gehan et al. 2018). Aside from a few exceptions, the spin parameters of the modes of all these types of pulsators in the co-rotating frame, , put them far into the super-inertial regime (completely to the right in Fig. 3). Almost all of these pulsators are hence slow rotators in our definition, with their forward modelling being done from their mode frequencies and the subsequent derivation of from Eq. (14) relying on the best forward seismic model. With the notable exception of the high-order p modes in the Sun-like star Bootis modelled with inclusion of the centrifugal deformation by Suárez et al. (2010), summaries of asteroseismic modelling results for the slow rotators can be found in García & Ballot (2019), Hekker & Christensen-Dalsgaard (2017), Lynas-Gray (2021), and Giammichele et al. (2022) for low-mass dwarfs, red giants, sub-dwarfs, and white dwarfs, respectively. Homogeneous analyses for the sub-giant phase were somewhat lacking after the Kepler mission finished, but are now also well underway (Ong et al. 2021; Noll et al. 2021), in anticipation of TESS light curves for this evolutionary phase.
@@ 4.1.3 Detection of internal rotation and core magnetism in red giants
The CoRoT space mission revealed that red giants are non-radial pulsators (De Ridder et al. 2009). It was a lucky circumstance that the mission programme could not avoid red giants in the observing fields dedicated to exoplanet research. The mode periods roughly range from half an hour to half a day, while the rotation periods of red giants typically range from ten days to hundreds of days. As discussed in the previous section, this implies that rotation can be treated perturbatively following the Ledoux approximation and that modelling can be done from the zonal modes. The CoRoT discovery of non-radial oscillations in red giants allowed the application of age-dating for various areas of the Milky Way (Hekker et al. 2009; Miglio et al. 2009); major advances in galactic archaeology (e.g. Miglio et al. 2013; Chiappini et al. 2015; Montalbán et al. 2021) is one of the important spin-offs of asteroseismology, as indicated in Fig. 1.
@@
Following the theoretical predictions from Dupret et al. (2009) triggered by CoRoT, dipole mixed modes were discovered in CoRoT data of a sub-giant (Deheuvels et al. 2010) and in Kepler data of a red giant (Beck et al. 2011). Subsequently, such modes were found in a whole sample of red giants by Bedding et al. (2011). The latter breakthrough study led to the important capacity to discriminate between stars on the red giant branch (RGB) and in the first or secondary red clump, even though such stars share the same surface properties. This discriminating capacity was also nicely illustrated for red giants in open clusters (Stello et al. 2011) and relies on the fact that the mixed modes have g-mode character in the deep interior of the star, yet are of acoustic nature in the envelope. Figure 3, helps us understand the nature of these modes: the p- and g-mode cavities delineated by the and symbols come closer to each other as a star evolves from being a dwarf to a red giant. As shown by Dupret et al. (2009), this narrows the zone between and in such a way that the waves can tunnel through both cavities without becoming fully evanescent in the transition zone, thus creating dipole mixed modes that probe the entire star.
@@
Core rotation frequencies were detected from the splitting of mixed dipole modes in a few Kepler red giants after two years of uninterrupted photometric monitoring (Beck et al. 2012). The positive effect of the longer duration of the Kepler light curves beyond two years implied major progress in seismic precision (Hekker et al. 2012). It led to a revolution in the understanding of angular momentum transport from large samples of evolved stars initiated by Mosser et al. (2012), with refined analyses from rotation inversions for a few RGB pulsators by Deheuvels et al. (2012) and red clump giants by Deheuvels et al. (2015). Following the initial breakthroughs, the internal rotation of evolved stars became an industrialised observational science of major importance (e.g. Gehan et al. 2018; Li et al. 2024b).
@@
Meanwhile, space asteroseismology delivered the internal rotation for several thousands of evolved low- or intermediate-mass stars, notably sub-giants and their successors climbing up the red giant branch, as well as giants in the first and secondary red clumps, sub-dwarfs, and white dwarfs. A summary of the rotational properties of about 1200 slow-rotating field stars is available in the review paper by Aerts et al. (2019) and is not repeated here. Figure 4 in that paper, along with the updated figures with measured internal rotation rates until the end of 2019 in Aerts (2021), place the internal rotational properties of low- and intermediate-mass stars in a global evolutionary picture. This revealed that the angular momentum of the core of helium burning red giants is in agreement with the angular momentum of white dwarfs.
@@
The major new insight from asteroseismology regarding internal rotation across stellar evolution is that the level of radial-differential rotation of single dwarfs of low and intermediate mass is low during their longest phases of evolution, when they fuse hydrogen into helium and helium into carbon and oxygen. As discussed in Aerts et al. (2019), who summarised all results obtained until mid-2018, any explanation of this asteroseismic picture of stellar evolution requires more efficient transport of angular momentum during these phases than anticipated prior to Kepler space asteroseismology. Said differently, much stronger coupling between the convective core and envelope occurs during the two central burning stages of intermediate-mass stars, while strong radial-differential rotation with a factor up to 20 for the core-to-envelope rotation rates occurs during the RGB (Li et al. 2024b). New theories to address this too low angular momentum transport in models of stars with a convective core have been developed. One way to increase the angular momentum transport while such stars evolve is to include internal gravity waves in the models. The occurrence of such waves was already proposed as an efficient way to transport angular momentum in low-mass stars like the Sun by Charbonnel & Talon (2005); Rogers & Glatzmaier (2006) and in high-mass stars by Rogers et al. (2013).
@@
Another way to achieve higher angular momentum transport in models is by including magnetic effects. The existence of internal magnetic fields in red giants was already inferred by Fuller et al. (2015) as an interpretation for depressed dipole mixed modes found in Kepler data for a small fraction of red giants. The accompanying study by Cantiello et al. (2016) showed that a magnetic greenhouse effect may be operational in red giants, turning g modes trapped in the core into Alfvén waves. The mode damping requires internal magnetic field strengths above about G, as confirmed by the theoretical predictions in Loi & Papaloizou (2017). A test of this greenhouse scenario for earlier phases of stellar evolution was developed by Stello et al. (2016), namely about half of the F dwarfs should have a strong internal magnetic field during their main-sequence phase. So far, it has not been possible to test this hypothesis from a representative asteroseismic population of F dwarfs, but internal magnetic field strengths of the 37 best modelled Kepler Dor pulsators inferred by Aerts et al. (2021) are in agreement with the scenario by Fuller et al. (2015) keeping in mind the minimal field strengths deduced by Cantiello et al. (2016).
@@
Although the interpretation by Fuller et al. (2015) was contested by Mosser et al. (2017), the presence of internal magnetic fields in a fraction of red giants is no longer a hypothesis. An observational breakthrough was achieved by Li et al. (2022a) and by Deheuvels et al. (2023), who found evidence of the presence of such fields inside the cores of respectively 3 and 11 red giants from 4 yr Kepler light curves. Li et al. (2022a) detected the core magnetism from the occurrence of asymmetrical rotational splittings in the triplets of mixed dipole modes. They interpreted this asymmetry as being due to the small effects induced by the Lorentz and Coriolis forces on the modes. In their approximation, the terms caused by and the magnetic force in are considered small compared to those due to the gas pressure gradient and gravity in Eq. (1), as confirmed by their numerical computations. Elegant analytical and numerical work to interpret the observational findings in the adopted approximation was offered by Li et al. (2022a, see the Supplementary Material) and by Mathis & Bugnet (2023) for various magnetic field topologies, following earlier theoretical studies on the effect of magnetism on modes of red giants by Loi & Papaloizou (2018); Gomes & Lopes (2020); Loi & Papaloizou (2020); Loi (2020b, a, 2021); Bugnet et al. (2021); Bugnet (2022). Deheuvels et al. (2023), on the other hand, inferred the presence of a core magnetic field in 11 RGB stars from the effect of the Lorentz force on the frequencies of zonal dipole modes. These 11 red giants do not reveal asymmetries in their dipole mode splittings. Rather, the Lorentz force affects the dipole zonal mode frequencies of consecutive radial order. The effect depends on the strength of the magnetic field and allows us to deduce lower limits on the core field strengths. As such, Deheuvels et al. (2023) found fields of 40 kG to 610 kG for these 11 stars. One of them, KIC 6975038, has depressed dipole modes, lending further support to the original interpretation of a magnetic greenhouse effect by Fuller et al. (2015).
@@
Meanwhile, Li et al. (2023a) detected a core magnetic field from asymmetrical splittings in 13 of the red giants with detected dipole mixed mode triplets found in the sample of 8000 compiled by Yu et al. (2018) and Gehan et al. (2018). The way the authors selected these 1200 RGB stars disfavours such stars with depressed dipole modes, as well as stars in an advanced stage of evolution along the RGB. The average core magnetic field strengths of these 13 stars range from 20 to 150 kG, representing a range from 5% to 30% of the critical field strength above which magneto-gravity waves are dissipated in the core rather than being propagative (Fuller et al. 2015; Rui & Fuller 2023). The detected fields were found to have a variety of horizontal field geometries. The core magnetic fields detected in less than 2% of the Kepler red giants with dipole mixed modes by Li et al. (2023a) and Deheuvels et al. (2023) are important to improve stellar evolution models. The low fraction found in the large homogeneous database of the Kepler 4 yr light curves does not necessarily imply the absence of a field for the majority of stars as the fields may remain undetected for large parts of the evolution. The field strengths were found to decrease as the cores of the stars shrink along their evolutionary path. The magnetic red giants have core rotation properties fully in line with those of the thousands of red giants without detected magnetic signatures. Both findings suggest that the magnetic fields do not cause much extra transport of angular momentum compared to the case of a non-detectable magnetic field. The sample of red giants with detected core fields is currently too small to make strong general conclusions. The effects of the core fields on internal mixing, if any, remain to be studied given the recent character of these discoveries. This may help to assess whether the magnetic field has any effect on element transport as the red giants evolve.
@@ 4.2 Asteroseismology of supernova progenitors: Low-order p and g modes in Cep stars
Moving somewhat to the left on the frequency axis in Fig. 3, we find moderate rotators among the Cep stars, such as HD 192575 already discussed above. The Cep stars are dwarfs or (super)giants of spectral type O or early B with masses roughly in the range 8 to 25 M⊙. They are hence precursors of supernovae. This class of pulsators has always played a special role in asteroseismology, and this is no different in this review. As early as 1902, Frost (1902) discovered radial-velocity variations for the prototype of the class, the star Cephei. At that time, the only explanation was that the star is a spectroscopic binary. Although the variability of tens of similar stars was misinterpreted in terms of orbital motion for about half a century, a breakthrough was achieved thanks to the three periodicities found in the radial-velocity time series for the class member CMa (Struve 1950; Van Hoof & Struve 1953). These properties in fact led Ledoux (1951) to write his breakthrough paper on non-radial oscillation modes in a rotating star, introducing the concept of rotationally split multiplets due to the joint effect of the Coriolis force and the projection in the line of sight for an observer, as discussed in Sect. 3.3. Even though we know today that Ledoux’s interpretation of the variability of CMa was wrong (e.g. Mazumdar et al. 2006, for a modern asteroseismic interpretation), it led him to derive a general description of the observed characteristics of non-radial oscillations in a rotating star, which is the one still in use today for slow rotators with modes in the super-inertial regime.
@@
Following Goupil et al. (2000), we assess the relative importance of second-order effects due to the Coriolis force, which scales as , and the effects due to the centrifugal force, which is proportional to . Thus, the relative role of the centrifugal versus Coriolis forces in any second-order theory in behaves as
@@
This ratio shows that the Coriolis force dominates over the centrifugal force for high-order g modes in many of the brightest SPB pulsators for which this ratio has been measured, and is typically between 0.01 and 0.1 (see Fig. 17 in De Cat & Aerts 2002). One can therefore ignore the oblateness caused by the centrifugal force for such moderately rotating pulsators, but cannot ignore the Coriolis force in their asteroseismic modelling (Aerts et al. 2021). Ballot et al. (2010) unravelled the domains of validity of a perturbative approach for the rotation to compute modes in the super-inertial regime and concluded that the second-order approach still gives satisfactory results for equatorial velocities up to 100 km s-1 for Dor stars and up to 150 km s-1 for SPB stars. A third-order perturbative approach increases the domains of validity by a few tens of km s-1. However, we have known since the era of space asteroseismology that most of these pulsators have their modes in the sub-inertial regime (see Fig. 3 and Aerts et al. 2021) where perturbative approaches for the rotation are not valid. We treat such high-order g-mode pulsators by adopting the traditional approximation of rotation (TAR) in Sect. 4.4, where we explain that the TAR is a mathematically elegant approximation that ignores the horizontal component of the rotation vector.
@@
The ratios from Eq. (15) for typical p modes in Cep or Sct stars are roughly between 1 and 10 (Goupil et al. 2000). This explains why the super-inertial modes in slowly to moderately rotating Cep and Sct stars can be described by a second-order perturbative theory. Ballot et al. (2010) evaluated the validity domain for these p-mode pulsators and concluded that this approximation is fine for equatorial rotation velocities up to about 70 km s-1. They noted that adding third-order terms hardly improves this because p modes are only weakly sensitive to the Coriolis force. For the cases where low-order linear oscillation modes in moderately rotating Sct or Cep stars can still be described perturbatively, we refer to Soufi et al. (1998), Karami (2008), and Guo et al. (2024) for appropriate mode expressions.
@@
Rotationally split p and g modes should in principle be easy to recognise for the slower rotators in the class and help identify the modes. However, despite the hundreds of OB pulsators found in space photometry (Aerts et al. 2006a; Balona et al. 2011; Balona 2016; Balona et al. 2019; Pedersen et al. 2019; Burssens et al. 2019, 2020; Balona 2022), the frequency spectra turn out to contain so many frequencies that the modes are hard to identify from the single-passband photometric variations. In particular, zonal modes are hard to find. As a result, forward modelling based on the fitting of individual identified modes has been done for only a few of these supernova progenitors from space photometry (Aerts et al. 2006b; Handler et al. 2009; Briquet et al. 2011; Aerts et al. 2011; Daszyńska-Daszkiewicz et al. 2017; Handler et al. 2019). Given their mass range and extensive convective cores, the forward modelling problem to fit the zonal modes is at least five-dimensional because convective core overshooting cannot be ignored and is an uncalibrated mixing phenomenon requiring at least one fit parameter.
@@
Some of the most detailed forward models for this class result from earlier months-long ground-based multi-site campaigns for several bright class members (Pamyatnykh et al. 2004; Briquet et al. 2007; Dziembowski & Pamyatnykh 2008; Briquet et al. 2012). A detailed comparison among the forward models resulting for the three pulsators HD 129929, Oph, and V2052 Oph, which have very similar dominant modes including an , g1 quintuplet, a radial mode, and an , p1 triplet, revealed that the magnetic star V2052 Oph has lower core overshooting than the two non-magnetic pulsators. This hinted at inhibition of mixing by a magnetic field.
@@
To date, internal rotation profiles assuming shellular rotation have been determined for five Cep stars. This had led to ratios of the near-core and envelope rotation rates between 1 and 5 (see Table 1 and Fig. 5 in Burssens et al. 2023, for a summary), where the rigid rotation case occurs in the only close binary pulsator among this small sample. However, the methods adopted to derive the internal rotation are diverse. For the extensively studied star Eri, Dziembowski & Pamyatnykh (2008) adopted the first-order Ledoux approximation according to Eq. (14) and used the fundamental radial mode and the two triplets g1 and p1 to obtain a core-to-envelop rotation ratio of about 5. Suárez et al. (2009), however, performed a more detailed analysis based on a second-order perturbative approach, including the p2 triplet. They found a factor lower envelope than near-core rotation from the three observed triplets, while also being able to explain most (but not all) of the measured triplet asymmetries. The approach by Suárez et al. (2009) is a promising route to understand angular momentum redistribution and chemical mixing due to radial-differential rotation and its induced turbulence in moderately rotating Cep stars.